Using Gauss’s theorem, calculate the flux of the vector field F =xˆi + y ˆj + zkˆ
through the surface of a cylinder of radius A and height H, which has its axis
along the z-axis. The base of the cylinder is on the xy plane.
1
Expert's answer
2020-03-17T10:01:58-0400


For a vector field


"\textbf{F}=P\textbf{i}+Q\textbf{j}+R\textbf{k}=x\textbf{i}+y\textbf{j}+z\textbf{k},"


the flux through the 3-dimensional surface element "S" is


"\Phi=\iint\textbf{F}\text{ d}S=\iiint\text{ div}\textbf{F}\text{ d}V,\\
\space\\
\text{ div}\textbf{F}=\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}+\frac{\partial F_z}{\partial z}=\\
\space\\
=\frac{\partial x}{\partial x}+\frac{\partial y}{\partial y}+\frac{\partial z}{\partial z}=\\
\space\\
=1+1+1=3.\\
\space\\
\iiint\text{ div}\textbf{F}\text{ d}V=\iiint3\text{d}V=3V=3\pi A^2H."